# Routing in Quantum Repeater Networks with Mixed Efficiency Figures

### v3

@@ III-A System Design
where $\eta$ is the efficiency of the qubit measurement in the basis states. Consider a scenario where the correct measurement basis for a qubit is $|1\rangle\langle 1|$. Then, during entanglement swapping in quantum repeaters, the measurement is characterized by the Eq. (2). For an ideal measurement device in quantum repeater $\eta=1$, the measurement results are entirely reliable. However, if $10\leq\eta<$ 1, there exists a non-zero probability of $1-\eta$ that the results will be incorrect. The same probability of error applies to measurement in the $|0\rangle\langle 0|$ basis described by Eq. (1).
@@ IV-A Simulation tool and Methodology
We constructed a specialized Python simulator for the system outlined in Section III-A. The code-base for all simulations is freely available on GitHub in the repository $vk9696/quantum-routing1$. This simulator models a quantum network that follows a time-slotted model, analogous to methodologies described in prior research [13, 14, 15].
@@ IV-B4 Blocking Probability Study
In the comparative analysis depicted in Figure 6, it is evident that the efficiency-aware approach yields a markedly lower blocking probability than shortest-path approach when the fidelity threshold, $\bar{F}$, is set at ${0.53,0.7}$. As $\bar{F}$ approaches 0.8, however, the disparity between the two methods diminishes. This convergence in performance can be attributed to the high fidelity threshold nearing 0.8, at which point the blocking probability approaches unity. Consequently, at such elevated thresholds, the opportunity for the efficiency-aware approach to further improve performance becomes negligible. In conclusion, the efficiency awareness technique improves the network performance in terms of blocking probabilities, i.e., fewer paths would be blocked upon using the efficiency awareness technique.

### v4

@@ Abstract
Note. This is the author accepted manuscript.
Please cite as: V. Kumar, C. Cicconetti, M. Conti and A. Passarella, “Routing in Quantum Repeater Networks with Mixed Efficiency Figures,” 2024 IEEE Future Networks World Forum (FNWF), Dubai, United Arab Emirates, 2024, pp. 198–203. DOI: 10.1109/FNWF63303.2024.11028804
@@ III-A System Design
where $\eta$ is the efficiency of the qubit measurement in the basis states. Consider a scenario where the correct measurement basis for a qubit is $|1\rangle\langle 1|$. Then, during entanglement swapping in quantum repeaters, the measurement is characterized by the Eq. (2). For an ideal measurement device in quantum repeater $\eta=1$, the measurement results are entirely reliable. However, if $0\leq\eta<1$, there exists a non-zero probability of $1-\eta$ that the results will be incorrect. The same probability of error applies to measurement in the $|0\rangle\langle 0|$ basis described by Eq. (1).
@@ IV-A Simulation tool and Methodology
We constructed a specialized Python simulator for the system outlined in Section III-A. The code-base for all simulations is freely available on GitHub in the repository vk9696/quantumrouting1vk9696/quantum-routing1. This simulator models a quantum network that follows a time-slotted model, analogous to methodologies described in prior research [13, 14, 15].
@@ IV-B4 Blocking Probability Study
In the comparative analysis depicted in Figure 6, it is evident that the efficiency-aware approach yields a markedly lower blocking probability than shortest-path approach when the fidelity threshold, $\bar{F}$, is set at 0.53,0.7{0.53,0.7}. As $\bar{F}$ approaches 0.8, however, the disparity between the two methods diminishes. This convergence in performance can be attributed to the high fidelity threshold nearing 0.8, at which point the blocking probability approaches unity. Consequently, at such elevated thresholds, the opportunity for the efficiency-aware approach to further improve performance becomes negligible. In conclusion, the efficiency awareness technique improves the network performance in terms of blocking probabilities, i.e., fewer paths would be blocked upon using the efficiency awareness technique.